# Getting to the Bottom of Serverless Billing

### v1

@@ 5. Related Work
(1) External characterization of serverless systems: Numerous studies characterized serverless systems from users’ perspective. Some investigated billing practices and cost efficiency of serverless platforms (Adzic and Chatley, 2017; Baldini et al., 2017; Eivy and Weinman, 2017; Lee et al., 2018; Yu et al., 2020; Liu and Niu, 2023; Cvetković et al., 2023). However, none offered a holistic top-down analysis like our study on how today’s serverless billing practices, request patterns, architectural overheads, and OS scheduling translate to inflated user costs. Other works performed cross-platform characterizations of resource allocation patterns and performance variations of serverless offerings (Wang et al., 2018; Kim and Lee, 2019; Yu et al., 2020; Copik et al., 2021; Wen et al., 2025b, a). They did not capture some critical factors that greatly impact serverless costs, such as concurrency models, resource allocation behaviors during KA, and OS scheduling granularity. Moreover, some of their measurements, such as default keep-alive durations and serving architectures, are now outdated due to the rapid evolution of serverless platforms.
(2) Characterization of production serverless workloads: Several provider-led studies characterized serverless workloads running within their systems (Shahrad et al., 2020b; Wang et al., 2021; Joosen et al., 2023, 2025; Zhang et al., 2021; Mahgoub et al., 2022). These provided valuable insights and enabled the trace-based analysis in parts of our study (e.g., quantifying inflated billable resources). However, none of these studies delivers an in-depth analysis that correlates the internal platform characteristics (e.g., overheads of architectures and scheduler settings) with the high costs experienced by serverless users. Our work fills this gap by holistically analyzing billing models and practices and measuring architectural overheads and OS scheduling effects, and connecting them to cost implications.
(3) Open-source serverless solutions, such as Knative (Knative, 2025), AWS Runtime API (AWS, 2025), Azure Functions Host (Azure, 2025), and workerd (Cloudflare, 2025b), enabled us to demystify overheads hidden in modern serverless serving infrastructures, which have not been reported before.
@@ 6. Conclusion
For the first time, we holistically demystify the high cost of serverless by conducting a comprehensive characterization of driving factors in a top-down manner from user-facing billing models, through architectural overheads, and finally to OS scheduling. We provide novel insights into how current billing practices, request serving architectures, KA resource allocation behaviors, and OS scheduling granularity contribute to inflated costs of serverless. These insights spark future directions for serverless practitioners on optimizing the cost-efficiency of serverless systems.

### v2

@@ 6. Related Work
(1) External characterization of serverless systems: Numerous studies characterized serverless systems from the users’ perspective. Some investigated billing practices and cost efficiency of serverless platforms (Adzic and Chatley, 2017; Baldini et al., 2017; Eivy and Weinman, 2017; Lee et al., 2018; Yu et al., 2020; Liu and Niu, 2023; Cvetković et al., 2023). However, none offered a holistic top-down analysis like our study on how today’s serverless billing practices, request patterns, architectural overheads, and OS scheduling translate to inflated user costs. Other works performed cross-platform characterizations of resource allocation patterns and performance variations of serverless offerings (Wang et al., 2018; Kim and Lee, 2019; Yu et al., 2020; Copik et al., 2021; Wen et al., 2025b, a). They did not capture some critical factors that greatly impact serverless costs, such as concurrency models, details of resource allocation during keep-alive, and OS scheduling granularity. Moreover, some of their measurements, such as default keep-alive durations and serving architectures, are now outdated due to the rapid evolution of serverless platforms.
(2) Characterization of production serverless workloads: Several provider-led studies characterized serverless workloads running within their systems (Shahrad et al., 2020b; Wang et al., 2021; Joosen et al., 2023, 2025; Zhang et al., 2021; Mahgoub et al., 2022b). These provided valuable insights and enabled the trace-based analysis in parts of our study (e.g., quantifying inflated billable resources). However, none of these studies delivers an in-depth analysis that correlates the internal platform characteristics (e.g., overheads of architectures and scheduler settings) with the high costs experienced by serverless users. Our work fills this gap by holistically analyzing billing models and practices and measuring architectural overheads and OS scheduling effects, and connecting them to cost implications.
(3) Open-source serverless solutions, such as Knative (Knative, 2025b), AWS Runtime API (Services, 2025b), workerd (Cloudflare, 2025e), and Azure Functions Host (Azure, 2025d), enabled us to demystify overheads hidden in modern serverless serving infrastructures, which have not been reported before.
@@ 7. Conclusion
For the first time, we holistically demystify the high cost of serverless by conducting a comprehensive characterization of driving factors in a top-down manner from user-facing billing models, through architectural overheads, and finally to OS scheduling. We provide novel insights into how current billing practices, request serving architectures, keep-alive resource allocation behaviors, and OS scheduling granularity contribute to the inflated costs of serverless. These insights spark future directions for serverless practitioners on optimizing the cost-efficiency of serverless systems.
Acknowledgements.
We thank the anonymous reviewers, and specially our shepherd, Andrea Arpaci-Dusseau, for helping us improve the paper. We also thank Alain Tchana for his feedback. This work was supported by the Natural Sciences and Engineering Research Council of Canada (NSERC), through research grants RGPIN-2021-03714 and DGECR-2021-00462, the Canada Graduate Scholarship (CGS D) program, and the Undergraduate Student Research Awards (USRA) program. This work was also supported in part by the Institute for Computing, Information and Cognitive Systems (ICICS) at UBC. Cloud resources from the Digital Research Alliance of Canada (RAS and RAC allocations) facilitated our research.